OpenStudy (anonymous):

4 3^sqrt 11 -7 4^sqrt11 +6 3^sqrt11 +13 4^sqrt 11

OpenStudy (anonymous):

@genius12

OpenStudy (anonymous):

10 3^ sqrt 11 + 6 4^ sqrt11 ?

OpenStudy (espex):

Yes, good job.

OpenStudy (anonymous):

multiply (3 sqrt5-1)(6sqrt5-4) @eSpeX

OpenStudy (anonymous):

@genius12

OpenStudy (anonymous):

4 + 18 sqrt 5 ???

OpenStudy (espex):

Is that \((3\sqrt{5}-1)(6\sqrt{5}-4)\)?

OpenStudy (anonymous):

yes

OpenStudy (espex):

You will FOIL it as you would any other set of factors, so First, then Outer, then Inner, then Last.

OpenStudy (anonymous):

Use FOIL:\[\bf (3\sqrt5-1)(6\sqrt5-4)=18(5)-18\sqrt{5}+4=..?\]Finish the simplifcation.

OpenStudy (espex):

F = \(3\sqrt{5} * 6\sqrt{5}\) etc..

OpenStudy (anonymous):

94sqrt

OpenStudy (anonymous):

5

OpenStudy (anonymous):

@eSpeX